### v2

# It’s Not You, It’s Clipping: A Soft Trust-Region via Probability Smoothing for LLM RL

Abstract
Training large language models (LLMs) with reinforcement learning (RL) methods such as PPO and GRPO commonly relies on ratio clipping to stabilise updates. While effective at preventing instability, clipping discards information, introduces gradient discontinuities and can prevent exploration of better policies. Inspired by label smoothing, we propose Probability Smoothing Policy Optimisation (PSPO). PSPO smooths current policy probabilities toward the behaviour policy before computing importance ratios, creating a soft trust region that preserves gradients while preventing destabilising updates. Unlike prior soft clipping approaches that use sigmoid-based transformations which can suffer from vanishing gradients and saturation, our method uses a linear interpolation, providing simpler and more robust gradient preservation.
Empirically, GR-PSPO outperforms clipping and sigmoid-based alternatives on mathematical reasoning benchmarks when refining models with prior domain knowledge, achieving an accuracy of 79.9% on GSM8K and 59.6% on MATH for Qwen2-Math-1.5B.
1 Introduction
Reinforcement learning (RL) is now a central component of large language model (LLM) fine-tuning. RL is commonly used after supervised fine-tuning (SFT) (Ouyang et al., 2022), where the aim is to refine an existing policy whilst preserving prior knowledge. In this setting, the main challenge is to make improvements without degrading existing knowledge, as unstable, destructive updates can waste expensive training. RL post-training has shown enhanced reasoning and can maintain context in long and complex conversations. This is especially important with the prevalence of Small Language Models, that due to their size are energy efficient and easier to control in a workflow as they can be trained and maintained by smaller companies but where these models in particular struggle to provide outputs that are understandable by a human user.
To stabilise updates, policy optimisation methods rely on constraints that limit deviation from a reference policy. In practice, this requires a balance between learning speed and stability, as overly conservative updates can make learning inefficient or infeasible. Trust Region Policy Optimisation (TRPO) (Schulman et al., 2015) constrains updates using the KL divergence, which allows for larger steps but is still computationally inefficient. Proximal Policy Optimization (PPO) (Schulman et al., 2017) provides an empirically stronger regularisation by using clipped probability ratios as a first-order approximation of the KL divergence. PPO clips the importance sampling ratio to effectively reduce the step size when the policy is outside the trust region, and underpins systems such as WebGPT (Nakano et al., 2021), LLaMA-2 Chat (Touvron et al., 2023), and Sparrow (Glaese et al., 2022).
This is extended in Group Relative Policy Optimization (GRPO) which is an adaptation of PPO specifically for LLMs  (Shao et al., 2024), inheriting the clipped probability, and has been applied to mathematical reasoning tasks (Shao et al., 2024), alongside other RL approaches (Luong et al., 2024; Mitra and Ulukus, 2025; Luo et al., 2025; Zheng et al., 2025). Some implementations of GRPO (Hugging Face, ) default to using a single pass over data, effectively avoiding clipping by only using on-policy data. This leads to the importance sampling ratio always being 1, which essentially reverts the approach back to a vanilla policy gradient method, and as such, will typically require small steps and be sample inefficient.
However, ratio clipping has drawbacks, namely vanishing gradients when the policy ratio leaves the clip range. Additionally, clipping can miss better policies outside of the clipped policy space (Chen et al., 2023), especially in problems where greater exploration might be beneficial. Alternatives to clipping have therefore emerged, predominantly focusing on smooth transforms to remove the brittleness of hard clipping. P3O (Chen et al., 2023) introduced sigmoid-based soft clipping through its SCOPIC objective to enable exploration beyond the trust region defined by typical hard clip boundaries. P3O was empirically shown to perform better than PPO on five deep RL gym environments, but has not been applied to GRPO or used for LLM fine-tuning prior to this work. Sigmoid methods can, however, lead to saturation of values at extreme ratios, reintroducing vanishing gradients in precisely the regions where clipping also fails. Recent work has addressed the asymmetric behaviour of positive versus negative advantages in LLM fine-tuning specifically, Tapered Off-Policy REINFORCE (TOPR) (Roux and Bellemare, 2025) employs asymmetric importance sampling to differentially weight positive and negative advantages showing better performance on GSM8K compared with PPO, DPO and REINFORCE on 8B parameter models. Soft Adaptive Policy Optimization (SAPO) (Gao et al., 2025) combines sigmoid soft gating with asymmetric control, similar to ideas from both P3O and TOPR, showing an increase in performance compared with GRPO and GSPO on mathematical tasks on 30B parameter models. Asymmetric methods require at least one additional parameter to be tuned, and TOPR’s performance seems to be quite dependent on the ratio of positive to negative samples in the training, which is not always known a priori or easy to control especially given that it will change across training. Other approaches include KL early stopping (Sun et al., 2022), though these can be brittle or saturating, particularly in complex settings.
We propose Probability Smoothing Policy Optimisation (PSPO), as an alternative to clipping. Instead of truncating ratios, we smooth the current policy’s probabilities toward the old behaviour policy before computing the importance ratio. This is inspired by label smoothing in supervised learning. This smoothing reduces overconfidence in any single action while retaining informative gradients everywhere.
We empirically evaluate PSPO by instantiating it within GRPO (Shao et al., 2024) (GR-PSPO) across three model types: Qwen2.5-0.5B, Qwen2.5-1.5B, and Qwen2-Math-1.5B. This experimental design allows us to test our hypothesis that PSPO’s advantages emerge when stability is required, such as with models with prior domain knowledge (the Math model).
2 Probability Smoothing Policy Optimisation
Policy gradient methods optimise the expected reward by updating the policy $\pi_{\theta}$ with respect to sampled trajectories. To effectively reuse trajectories from an old policy $\pi_{\theta_{\text{old}}}$, any update is regularised using importance sampling. Importance sampling estimates how likely the (state $s_{t}$, action $a_{t}$) pair would occur given the current policy. In PPO (Schulman et al., 2017) and GRPO Shao et al. (2024), this is approximated with a ratio of the current policy $\pi_{\theta}$ and the old, behaviour policy $\pi_{\theta_{\text{old}}}$ which generated the trajectory. This ratio is defined in equation  1,
$r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}.$ (1)
GRPO is an adaptation of PPO,  (Shao et al., 2024), which removes the need for a critic model, reducing the amount of training resources and developed specifically for LLMs. GRPO samples a group $G$ of outputs $a$ for a given prompt $s\in S$, and uses the group scores $r$ as a baseline estimate to then calculate the advantage $\hat{A}$ using the relative rewards based on the current group baseline; $\hat{A}_{t,i}=R_{t,i}-\bar{R}_{t,i}$. GRPO includes the same clipping principle as PPO in its surrogate objective, although some default implementations suggest that using GRPO with only 1 iteration over the data gives comparable performance and negates the effect of clipping. GRPO aims to maximise the objective function:
$\begin{split}J^{\text{GRPO}}(\theta)=\mathbb{E}_{t}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Bigg\{\min\Big(r_{t,i}(\theta)\hat{A}_{t,i},\,\text{clip}(r_{t,i}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{t,i}\Big)-\beta\mathbb{D}_{KL}[\pi_{\theta}\mid\pi_{ref}]\Bigg\}\Bigg],\end{split}$ (2)
where $\mathbb{D}_{KL}[\pi_{\theta}\mid\pi_{ref}]$ is an estimate of the KL divergence from the current policy $\pi_{\theta}$ to a reference policy $\pi_{ref}$, and $\beta$ is a hyper-parameter which controls the strength of this penalty. This KL divergence is similar to that used by TRPO, but in GRPO it is used as a soft penalty of $\pi_{\theta}$ to $\pi_{ref}$ compared with TRPO’s hard constraint of $\pi_{\theta_{\text{old}}}$ to $\pi_{\theta}$. In some popular implementations (Hugging Face, ), $\beta$ is set to 0 as it reduces memory usage and improves the training speed by not needing to load the reference model.
In complex RL problems, there is often multiple optimal actions. Language generation tasks demonstrate this excellently, as within language, there are typically many possible words (actions) that can represent the same meaning (achieve the same goal).
To reduce overconfidence in any single action in a given state, we took inspiration from the label smoothing regularisation method used in supervised learning (Szegedy et al., 2016). Label smoothing has been shown to reduce overconfidence and improve the robustness of a model (Müller et al., 2019) (Goibert and Dohmatob, 2019). Label smoothing, equation  3, moves from one-hot encoded target distribution $\varphi(k\mid x)$ to soft targets $\tilde{\varphi}(k\mid x)$ that are a weighted average of the hard target distribution and another distribution, traditionally the uniform distribution $u(k)$ (Szegedy et al., 2016),  [^1] In the original label smoothing paper, $\varepsilon$ is used as the smoothing parameter, we use $\alpha$ to avoid confusion with the clipping range, often denoted as $\varepsilon$.
$\tilde{\varphi}(k\mid x)=(1-\alpha)\cdot\varphi(k\mid x)+\alpha\cdot u(k),$ (3)
where $\alpha\in[0,1]$ controls the smoothing strength.
We apply (3) to the current policy probability in equation  4,
$\tilde{\pi_{\theta}}(a_{t}\mid s_{t})=(1-\alpha)\pi_{\theta}(a_{t}\mid s_{t})+\alpha\cdot q(a_{t}\mid s_{t}),$ (4)
where $q(\cdot)$ represents the distribution we want to smooth towards.
For policy optimisation, updates should be within a trust region to enable larger step updates. Therefore, we decided to smooth towards the old behaviour policy, $q=\pi_{\theta_{\text{old}}}$, rather than the uniform distribution, so the smoothing behaves as a behaviour-anchored trust region.  Szegedy et al. (2016) noted that the change in the loss when using label smoothing could be equivalently considered as an estimate of the KL divergence. Given that PPO used the probability ratio clipping as a first-order estimate of KL-divergence for regularisation, this bolsters our intuition to smooth towards the old policy, and by doing so, we introduce an equivalent estimate of the KL-divergence with the smoothed action probability and create a soft trust region. The smoothed probability becomes equation  5,
$\tilde{\pi_{\theta}}(a_{t}\mid s_{t})=(1-\alpha)\pi_{\theta}(a_{t}\mid s_{t})+\alpha\cdot\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t}).$ (5)
If we then consider the ratio equation, we can find the effect from the smoothed probability on the ratio using equation  5,
$\tilde{r}_{t}(\theta)=\frac{\tilde{\pi_{\theta}}(a_{t}\mid s_{t})}{{\pi_{\theta_{\text{old}}}}(a_{t}\mid s_{t})},$ (6)
which given (5), becomes equation 7,
$\begin{split}\tilde{r}_{t}(\theta)=\frac{(1-\alpha)\pi_{\theta}(a_{t}\mid s_{t})+\alpha\cdot\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}\\ =(1-\alpha)r_{t}+\alpha.\end{split}$ (7)
Having derived the smoothed ratio (Eq. 7) we now establish formally that this simple linear transformation provides the properties we intuitively assumed: (i) contraction toward the behaviour policy in both total variation and KL divergence, ensuring updates remain within a trust region; (ii) preservation of non-zero gradients everywhere, avoiding the learning signal loss caused by clipping; and (iii) regularisation against overconfidence in any single action. We state these results as formal propositions.
Soft Trust Region - Implicit divergence control from probability smoothing.
Given the smoothed policy and ratio (Eqs. ( 4)–( 7)), the linear interpolation, $\tilde{r}_{t}(\theta)=(1-\alpha)r_{t}+\alpha$, yields a contraction around $r=1$ and induces a soft trust region anchored at $\pi_{\theta_{\text{old}}}$, consistent with our intuition.
Lemma 2.1 (Total variation (TV) contraction).
For any state $s$ and $\alpha\in[0,1]$,
$\|\tilde{\pi}_{\theta}(\cdot\mid s)-\pi_{\theta_{\text{old}}}(\cdot\mid s)\|_{1}=(1-\alpha)\,\|\pi_{\theta}(\cdot\mid s)-\pi_{\theta_{\text{old}}}(\cdot\mid s)\|_{1}.$
Proof.
Since $\tilde{\pi}_{\theta}-\pi_{\theta_{\text{old}}}=(1-\alpha)(\pi_{\theta}-\pi_{\theta_{\text{old}}})$ pointwise, linearity of the $\ell_{1}$ norm gives the result directly. ∎
This means $\alpha$ directly controls the maximum distance the smoothed policy can drift from the behaviour policy: setting $\alpha=0.1$ guarantees the smoothed policy is at most 90% as far from $\pi_{\theta_{\text{old}}}$ as the unsmoothed policy would be, in total variation distance. This provides a tunable constraint on policy updates without hard truncation.
Corollary 2.2 (KL upper bounds shrink under smoothing).
We use the joint convexity of KL and set $\lambda=1-\alpha$, $P_{1}=\pi_{\theta}$, $P_{2}=\pi_{\theta_{\text{old}}}$, $Q_{1}=\pi_{\theta_{\text{old}}}$, $Q_{2}=\pi_{\theta_{\text{old}}}$. This gives us:
$\begin{split}\lambda P_{1}+(1-\lambda)P_{2}=(1-\alpha)\pi_{\theta}+\alpha\pi_{\theta_{\text{old}}}=\tilde{\pi_{\theta}},\\ \lambda Q_{1}+(1-\lambda)Q_{2}=(1-\alpha)\pi_{\theta_{\text{old}}}+\alpha\pi_{\theta_{\text{old}}}=\pi_{\theta_{\text{old}}}\end{split}$
Given that $D_{\mathrm{KL}}\!\big(\pi_{\theta_{\text{old}}}\|\ \pi_{\theta_{\text{old}}}\big)=0$, we then find:
$D_{\mathrm{KL}}\!\big(\tilde{\pi}_{\theta}\,\|\,\pi_{\theta_{\text{old}}}\big)\;\leq\;(1-\alpha)\,D_{\mathrm{KL}}\!\big(\pi_{\theta}\,\|\,\pi_{\theta_{\text{old}}}\big)$
Similarly for the reverse direction we find:
$D_{\mathrm{KL}}\!\big(\pi_{\theta_{\text{old}}}\,\|\,\tilde{\pi}_{\theta}\big)\;\leq\;(1-\alpha)\,D_{\mathrm{KL}}\!\big(\pi_{\theta_{\text{old}}}\,\|\,\pi_{\theta}\big).$
Hence $\alpha$ directly sets a soft trust-region radius in both total variation (TV) and (upper-bounded) KL.
This establishes that probability smoothing induces an implicit KL penalty, which can be viewed as a soft trust-region with radius of $\alpha$.
Proposition 2.3 (Ratio contraction and non-vanishing slopes).
For any action $a$ with $\pi_{\theta_{\text{old}}}(a\mid s)>0$, and $r(a)$ is the importance sampling ratio for action $a$,
$|\tilde{r}(a)-1|\leq(1-\alpha)\,|r(a)-1|,\qquad\frac{\partial}{\partial r}\big(\tilde{r}\,A\big)=(1-\alpha)A.$
Thus, PSPO preserves slope $(1-\alpha)A$ everywhere, avoiding the flat plateaus introduced by clipping outside $1-\varepsilon,1+\varepsilon]$ (Fig. 1).
This is the key advantage over clipping. With PPO/GRPO clipping, when $r>1+\varepsilon$ (for positive advantages) or $r<1-\varepsilon$ (for negative advantages), the gradient is exactly zero, the model receives no learning signal for these samples. PSPO instead scales the gradient by $(1-\alpha)$, which is non-zero everywhere. This means PSPO continues to learn from samples that clipping would ignore, while the contraction still prevents destabilising large updates.
Proposition 2.4 (Overconfidence regularisation).
For any state $s$ and action $a$, the smoothed policy satisfies:
$\tilde{\pi_{\theta}}(a\mid s)\leq max\Big(\pi_{\theta}(a\mid s),\pi_{\theta_{\text{old}}}(a\mid s)\Big),$
with strict inequality whenever $\pi_{\theta}(a\mid s)\neq\pi_{\theta_{\text{old}}}(a\mid s)$ and $\pi_{\theta}(a\mid s)>\pi_{\theta_{\text{old}}}(a\mid s)$.
Proof.
From the definition $\tilde{\pi_{\theta}}(a\mid s)=(1-\alpha)\pi_{\theta}(a\mid s)+\alpha\pi_{\theta_{\text{old}}}(a\mid s)$. When $\pi_{\theta}(a\mid s)\geq\pi_{\theta_{\text{old}}}(a\mid s)$, we have $\tilde{\pi_{\theta}}(a\mid s)<\pi_{\theta}(a\mid s)$. When $\pi_{\theta}(a\mid s)<\pi_{\theta_{\text{old}}}(a\mid s)$, we have $\tilde{\pi_{\theta}}(a\mid s)<\pi_{\theta_{\text{old}}}(a\mid s).$
This property prevents the policy from over fitting and becoming overly confident in any single action. When the current policy assigns higher probability to an action than the behaviour policy did, smoothing pulls this probability down. This is particularly relevant for language generation, where multiple tokens may be equally valid continuations.
Proposition 2.5 (PSPO surrogate as a scaled policy gradient with implicit stability).
The per-state PSPO objective can be written
$\begin{split}\mathcal{J}_{\text{PSPO}}(\theta)=\mathbb{E}_{a\sim\pi_{\theta_{\text{old}}}}\![\tilde{r}(a)\,A(a)]\\ =(1-\alpha)\,\mathbb{E}_{a\sim\pi_{\theta}}\![A(a)]+\alpha\,\mathbb{E}_{a\sim\pi_{\theta_{\text{old}}}}\![A(a)].\end{split}$
using equation 1 and the change of measure formula. Only the first term depends on $\theta$, and using policy gradient theorem, $\nabla_{\theta}\mathcal{J}_{\text{PSPO}{}}=(1-\alpha)\,\mathbb{E}_{a\sim\pi_{\theta}}\!\big[\nabla_{\theta}\log\pi_{\theta}(a\mid s)\,A(a)\big].$ Hence, PSPO is the on-policy gradient scaled by $(1-\alpha)$, while the policy itself is mixed with $\pi_{\theta_{\text{old}}}$ (Lemma 2.1), implicitly controlling divergence without an explicit KL term (which we set $\beta{=}0$ in our GRPO runs).
This shows that PSPO can be seen as performing standard on-policy gradient descent, but with two modifications: the gradient is scaled by $(1-\alpha)$, slowing learning, while the policy used to compute advantages is mixed with $\pi_{\theta_{\text{old}}}$. Together, these effects stabilise learning by anchoring updates to recent behaviour.
Proposition 2.5 shows that PSPO scales the policy gradient by a factor of $1-\alpha$. While this may resemble a simple reduction in the learning rate, PSPO is not equivalent to learning-rate scaling alone. The scaling arises from probability mixing with the behaviour policy, which simultaneously contracts the policy distribution toward $\pi_{\theta_{\text{old}}}{}$. As a result, PSPO adaptively alters the update magnitudes based on the model’s own prior probabilities, rather than applying a uniform step-size reduction. In contrast, clipping also implicitly rescales gradients, but does so discontinuously by truncating updates past a fixed ratio threshold.
Figure 1 illustrates how PSPO and clipping affect the ratio and the gradients. Clipping creates flat regions where gradients vanish: for $A>0$, the clipped surrogate is constant for $r>1+\varepsilon$; for $A<0$, it is constant for $r<1-\varepsilon$. In contrast, our method smooths the current policy toward the behaviour policy, giving us Eq. 7. This smooths ratios toward $1$, creating a soft trust region anchored by $\pi_{\theta_{\text{old}}}$, while maintaining non-zero gradients everywhere: $\tfrac{\partial}{\partial r}(\tilde{r}A)=(1-\alpha)A$. Therefore, PSPO preserves learning signal outside the clip range whilst still controlling updates.
Figure 1: Illustrative plot of ratio $r$ vs. the surrogate term AA, and the gradients for $A>0$ and $A<0$, with $\varepsilon=0.2$ and $\alpha=0.2$. For $A>0$ the clipped ratio is flat (zero gradient) for $r>1+\varepsilon$; for $A<0$, the clipped ratio is flat when $r<1-\varepsilon$. PSPO’s slope is $(1-\alpha)A$ everywhere, creating a soft trust region without hard plateaus.
Applicability.
PSPO is a direct replacement for ratio clipping, requiring only the substitution $\tilde{r}_{t}=(1-\alpha)r_{t}+\alpha$ for $r_{t}$ in any clipped-ratio objective. This change requires no additional computation or memory beyond evaluating the usual importance ratio.
Selecting $\alpha$.
The smoothing parameter $\alpha\in[0,1]$ controls the strength of the trust region by determining the fraction of the old policy to mix with the current policy. The larger the value of $\alpha$, the stronger the regularisation is, and the ratio is contracted more toward 1, and the KL upper bounds shrink (Corollary 2.2). A larger $\alpha$ is analogous to a smaller clipping range $\varepsilon$ in traditional GRPO. In practice, we have found that $\alpha\in[0.1,0.4]$ works well across our experiments, comparable to typical clipping ranges used in PPO/GRPO. Importantly, $\alpha$ replaces $\varepsilon$ as the trust region hyperparameter: both methods require tuning exactly one scalar controlling the constraint strength, so PSPO introduces no additional hyperparameter burden.
Application to GRPO:
We demonstrate how PSPO can apply to GRPO, denoted as GR-PSPO which changes (2) to equation  8,
$J^{\text{GR-PSPO}}(\theta)=\mathbb{E}_{t}\Bigg[\frac{1}{G}\sum_{i=1}^{G}(\tilde{r}_{t,i}(\theta)\hat{A}_{t,i})-\beta\cdot\mathbb{D}_{KL}[\pi_{\theta}\mid\pi_{ref}]\Bigg].$ (8)
3 Experimental Setup
Model and Prompt Formatting We fine-tune the open source causal LMs Qwen2.5-0.5B and -1.5B (Qwen Team, 2024) and also the Qwen2-Math-1.5B model which has specialised domain knowledge of mathematics (Yang et al., 2024) using their own tokeniser. The primary difference between the Math and the base models is that the Math model is a Qwen2-1.5B model which is then pretrained on a mathematics-specific corpus using supervised fine-tuning.
We use these three models as they are openly available, allow us to test scale across different, smaller sizes and enables us to compare models with and without existing mathematics knowledge. All runs use identical tokenisation and prompt formatting. Each sample is formatted with a system instruction followed by the user problem text. We use the model’s native chat template via tokenizer.apply_chat_template(…, add_generation_prompt=True) to append the assistant header and ensure the model completes in the assistant role, consistent with how the models are trained to be used.
When decoding completions, we set max_completion_length=128 with a maximum prompt length of 512 tokens, giving 640 tokens overall. Token count in LLM fine-tuning is a big contributing factor on the compute required (approx. 4 times the token count memory required for training), so we chose 128 completion tokens to balance computational constraints whilst allowing room for the response to include reasoning steps. This is also the approximate length of the GSM8K gold answers. We do not enforce any additional stop strings beyond the template defined end of string token.
Training All methods are trained using the same effective batch size and decoding settings. We train each method across 3 seeds, for the same number of tokens for fair comparison of the different methods; the best-validation checkpoint model is used for evaluations.
For the base models (Qwen2.5-0.5B and -1.5B), which lack specialised knowledge, we trained for 150M tokens. For Qwen2-Math-1.5B, which has existing mathematical knowledge to build on, we train for 75M tokens. Models with existing domain knowledge typically require less training to refine.
[…]
Table 1: Trust region methods evaluated. All use GRPO’s group relative advantages.
Method Trust Region Mechanism Implementation
GRPO-clip IS ratio clipping TRL baseline
[…]
P3O Sigmoidal soft clipping (SCOPIC) Adapted to GRPO
TOPR Asymmetric advantage gating Adapted to GRPO
SAPO Asymmetric sigmoidal scaling Backported from TRL
GR-PSPO Probability smoothing Novel (ours)
Hyperparameter Tuning To tune the hyperparameters, we used a grid-search across learning rate, clipping range, smoothing alpha strength, and tau values. Memory-related and decoding parameters were fixed across all methods. More details on the tuning method and the parameters used for each method can be found in Appendix A.
Evaluation We evaluate each test set across temperatures $T\in\{0.0,0.2,0.4,0.6,0.8\}$ with top-p=1.0p{=}1.0. We report the mean zero-shot Top-1 accuracy with 95% confidence intervals. We use Top-1 at $T=0$ as it reflects single-answer deployment settings and is deterministic. We calculate the accuracy using the true accuracy of a response, i.e. if anywhere in the response the correct answer appears. We also evaluate the reward function style accuracy as well.
3.1 Mathematical Reasoning Datasets
The base models were fine-tuned on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) (using the standard train/test split) and evaluated in-domain on the test sets and out-of-distribution on ASDiv (Miao et al., 2020) and SVAMP (Patel et al., 2021). These benchmarks span basic arithmetic word problems (ASDiv), robustness to linguistic perturbations (SVAMP), and competition-level reasoning (MATH; sampled from MATH (Hendrycks et al., 2021)). Following Minerva (Lewkowycz et al., 2022) and OpenWebMath (Paster et al., 2024), we restrict evaluation to problems with numeric final answers to enable automatic verification. For GSM8K training, we split the published training set into 7000 train and 472 validation examples.
3.1.1 System Prompt
System:
You are a careful math solver. Think through the solution and show the steps. Use English only. End the response with the final answer only in the format: ’#### <<final numeric answer only>>’.
User content: the raw problem text (with no few-shot exemplars).
We use this prompting to encourage stepwise reasoning and finish with a single, numeric final answer, which can be more easily extracted for calculating the reward.
3.1.2 Reward Function
Our rewards follow the commonly used correctness-based setup (Lewkowycz et al., 2022; Paster et al., 2024; DeepSeek-AI et al., 2025): R=1R{=}1 for exact numeric correctness within $10^{-6}$ tolerance, plus a +0.05+0.05 shaping bonus if the output matches the format “#### <<number>>”; values are constrained to $[0,1]$ giving $\{0,0.05,1\}$. We first attempt to extract the number from the requested format; if this is not present, we fall back to the last numeric token in the completion.
Table 2: Comparison of our results across all the model sizes on mathematical reasoning benchmarks. The accuracy presented is the true accuracy (i.e. has the correct answer anywhere in the response, even if it provides an incorrect answer at the end). Accuracy reported as average over 3 seeds, evaluated in greedy setting with T=0.0, and top_p=1.0.
Model Size Training Accuracy (%)
Data GSM8K ASDiv SVAMP MATH
Qwen2.5-0.5B 0.5B N/A 16.9 42.7 42.0 30.2
Qwen2.5-0.5B+GRPO-clip 0.5B GSM8K 54.1 77.8 66.6 43.0
Qwen2.5-0.5B+GRPO-noclip 0.5B GSM8K 56.0 77.9 70.0 42.6
Qwen2.5-0.5B+P3OChen et al. (2022) 0.5B GSM8K 44.5 68.5 58.3 38.9
Qwen2.5-0.5B+SAPOGao et al. (2025) 0.5B GSM8K 31.8 62.0 51.8 36.2
Qwen2.5-0.5B+GR-PSPO 0.5B GSM8K 53.0 78.8 69.7 41.9
Qwen2.5-0.5B+GRPO-clip 0.5B GSM8K+MATH 49.4 76.5 62.1 34.4
Qwen2.5-0.5B+GRPO-noclip 0.5B GSM8K+MATH 54.1 76.6 56.8 36.2
Qwen2.5-0.5B+P3OChen et al. (2022) 0.5B GSM8K+MATH 50.1 75.0 62.8 36.1
Qwen2.5-0.5B+SAPOGao et al. (2025) 0.5B GSM8K+MATH 50.3 76.6 63.3 34.5
Qwen2.5-0.5B+TOPRRoux and Bellemare (2025) 0.5B GSM8K+MATH 2.8 3.2 7.3 8.6
Qwen2.5-0.5B+GR-PSPO 0.5B GSM8K+MATH 50.9 75.9 65.7 37.5
Qwen2.5-1.5B 1.5B N/A 5.2 5.0 4.0 17.6
Qwen2.5-1.5B+GRPO-clip 1.5B GSM8K+MATH 73.7 88.8 83.3 45.9
Qwen2.5-1.5B+GRPO-noclip 1.5B GSM8K+MATH 75.4 87.9 83.7 48.0
Qwen2.5-1.5B+GR-PSPO 1.5B GSM8K+MATH 73.8 89.4 82.7 50.0
Qwen2-Math-1.5B 1.5B N/A 24.7 61.8 56.3 34.8
Qwen2-Math-1.5B+GRPO-noclip 1.5B GSM8K+MATH 78.4 92.8 84.6 58.3
Qwen2-Math-1.5B+GRPO-clip 1.5B GSM8K+MATH 70.3 91.6 82.0 52.3
Qwen2-Math-1.5B+P3OChen et al. (2023) 1.5B GSM8K+MATH 78.8 93.0 84.7 55.8
Qwen2-Math-1.5B+SAPOGao et al. (2025) 1.5B GSM8K+MATH 79.2 93.8 85.3 59.3
Qwen2-Math-1.5B+GR-PSPO 1.5B GSM8K+MATH 79.9 94.0 85.7 59.6
4 Performance of PSPO compared to the state-of-the-art
We evaluate GR-PSPO against multiple trust region baselines across mathematical reasoning tasks. Table 2 shows results across all model configurations. Our main finding is that GR-PSPO’s advantages grow with model size and domain specialisation, validating our hypothesis that probability smoothing is most effective when refining models that already have substantial knowledge, when stability is most important.
4.1 Model Capability and Trust Region Performance
GR-PSPO’s advantage scales with prior domain knowledge. On Qwen2.5-0.5B, differences between methods are small. On Qwen2.5-1.5B, GR-PSPO achieves the strongest MATH performance (+4.1pp over GRPO-clip). On Qwen2-Math-1.5B, a model with existing mathematical domain knowledge, GR-PSPO outperforms on all four benchmarks.
This scaling is reflected in GSM8K performance, where GR-PSPO transitions from -$\text{-}3.2$% relative to the best per dataset (0.5B) to -$\text{-}1.6$% (1.5B) to +0.7% (Math-1.5B).
Math-Specialised Model
On Qwen2-Math-1.5B, GR-PSPO achieves the strongest performance across all benchmarks. At $T=0$, GR-PSPO attains 79.9% on GSM8K (versus 78.4% for GRPO-noclip), 94.0% on ASDiv, 85.7% on SVAMP, and 59.6% on MATH.
The consistency across tasks—from arithmetic word problems (ASDiv, SVAMP) to competition-level reasoning (MATH)—suggests that probability smoothing generalises across difficulty levels when refining models with a base competence.
4.2 Base Models
On base Qwen2.5 models without prior mathematical training, GR-PSPO’s advantages are most pronounced on MATH, the most challenging benchmark. For Qwen2.5-0.5B trained on GSM8K+MATH, GR-PSPO achieves the highest MATH accuracy (37.5% vs. 36.2% for GRPO-noclip) but trails on GSM8K (50.9% vs. 54.1%). For Qwen2.5-1.5B, GR-PSPO achieves 50.0% on MATH versus 45.9% for GRPO-clip (+4.1pp), with comparable GSM8K performance.
When fine-tuning a base model on a new task, there is less existing knowledge to preserve, and aggressive updates may accelerate learning. When refining a model that already has domain expertise, stability becomes critical. GR-PSPO’s behaviour-anchored smoothing provides this property, explaining its consistent gains on Qwen2-Math.
4.3 Comparison to alternative methods.
GR-PSPO generally outperforms P3O and SAPO on both GSM8K and MATH across model sizes (Table 2). On Math-1.5B, GR-PSPO achieves 59.6% on MATH versus 55.8% for P3O (+3.8pp). We also adapted TOPR (Roux and Bellemare, 2025) to the GRPO framework, but the model seems to collapse during training, leading to below base model performance. TOPR was originally designed for REINFORCE, so the model may of collapsed due to the group-relative advantage formulation in GRPO.
4.4 Comparison to published results.
Our results outperform prior work at similar scales (Table 3). Luo et al. (2025) report 50.1% GSM8K and 21.2% MATH on WizardMath-Large (0.7B); our 0.5B GR-PSPO achieves 50.9% and 37.5% respectively. Our Qwen2-Math-1.5B results (79.9% GSM8K, 59.6% MATH) substantially exceed WizardMath-GPT2-XL (58.9%, 25.4%) at comparable scale. More notably, our Qwen2-Math-1.5B results outperform the Llama 3 8B model from Roux and Bellemare (2025), achieving +4.5pp on GSM8K and +36.9pp on MATH despite being substantially smaller. Direct comparison is complicated by differences in evaluation methodology; notably, our context window is much shorter (640 tokens with 128 completion tokens versus 2000+ tokens in prior work, or 128 completion tokens vs. the 512 completion tokens). We evaluate on the harshest settings, and still have competitive performance.
Table 3: Comparison on mathematical reasoning benchmarks. Reward Source: Rule = rule-based correctness function, Model = trained reward model. Reward Type: O = outcome (final answer), P = process (intermediate reasoning steps). The accuracy presented is the true accuracy (i.e. has the correct answer anywhere in the response, even if it provides an incorrect answer at the end). Note: “—” indicates data could not be found; “N/A” indicates not applicable.
Model Size Training Reward Evaluation Accuracy (%)
Data Method Source Type Prompting Decoding Tokens GSM8K MATH
Smaller Language Models << 1B parameters.
WizardMath-Small (Luo et al., 2025) 0.1B GSM8K+MATH PPO Model P+O - T=0 2048 26.4 12.3
WizardMath-Medium (Luo et al., 2025) 0.3B GSM8K+MATH PPO Model P+O - T=0 2048 38.7 15.6
WizardMath-Large (Luo et al., 2025) 0.7B GSM8K+MATH PPO Model P+O - T=0 2048 50.1 21.2
Qwen2.5-0.5B+RL (Zhuang et al., 2025) 0.5B GSM8K GRPO Rule O - T=0.6, top_p=0.95 32768 34.1 23.6
FlanT5-L Specialized (Fu et al., 2023) 0.76B GSM8K Distill N/A N/A Few-shot - 8192 20.2 -
GPT-2 (Socratic) (Shridhar et al., 2023) 0.77B GSM8K+SCoT Distill N/A N/A Socratic CoT - - 21.1 -
Qwen2.5-0.5B+GRPO-noclip 0.5B GSM8K+MATH GRPO (iterations=1) Rule O Zero-shot T=0, top_p=1.0 640 54.1 36.2
Qwen2.5-0.5B+GR-PSPO 0.5B GSM8K+MATH GR-PSPO Rule O Zero-shot T=0, top_p=1.0 640 50.9 37.5
Mid-to-Large Language Models >> 1B parameters.
GPT-2-XL (Brown et al., 2020) 1.5B GSM8K+MATH N/A N/A N/A Few-Shot - 2048 15.4 6.9
WizardMath-GPT2-XL (Luo et al., 2025) 1.5B GSM8K+MATH PPO Model P+O - T=0 2048 58.9 25.4
Llama 3 8B+TOPR Roux and Bellemare (2025) 8B GSM8K+MATH TOPR Rule O Few-Shot - 512 (completion tokens) 75.4 22.7
Qwen2.5-1.5B+GRPO-noclip 1.5B GSM8K+MATH GRPO (iterations=1) Rule O Zero-shot T=0, top_p=1.0 640 75.4 48.0
Qwen2.5-1.5B+GR-PSPO 1.5B GSM8K+MATH GR-PSPO Rule O Zero-shot T=0, top_p=1.0 640 73.8 50.0
Qwen2-Math-1.5B+GR-PSPO 1.5B GSM8K+MATH GR-PSPO Rule O Zero-shot T=0, top_p=1.0 640 79.9 59.6
Smoothing Alpha Sensitivity
We ablate $\alpha$ on Qwen2-Math-1.5B (Appendix B). Performance peaks around $\alpha$ = 0.4.
Preliminary larger model results (7B).
To further consider scaling behaviour we ran one exploratory seed for a reduced number of tokens on 7B models. The Qwen2.5-7B run shows a similar trend as our smaller-scale experiments, with: GR-PSPO outperforming on MATH. For example, PSPO’s top-1 (true) accuracy is 55.8% on MATH compared with the base model’s 33.9%, and GRPO-clip’s 53.6%. GRPO-noclip has an accuracy of 0% indicating complete collapse. These 7B measurements are single-seed and exploratory, so we do not use them to change conclusions drawn from the 3-seed experiments.
5 Discussion and Limitations
Unlike clipping which truncates the surrogate objective and discards gradient information outside the allowed range, PSPO contracts ratios toward 1 through probability smoothing while preserving non-zero gradients everywhere. This is achieved without additional compute, memory overhead, or extra optimisation objectives, the only change is a linear transformation of the importance ratio.
Our results across three model configurations validate that probability smoothing is most effective when stability is required, particularly when refining a model with existing domain knowledge.
This is shown by performance improvements scaling with domain specialisation. On general base models (Qwen2.5-0.5B/1.5B), GR-PSPO shows advantages primarily on the hardest benchmark (MATH). On the specialised model (Qwen2-Math-1.5B), GR-PSPO outperforms consistently across all tasks, suggesting that PSPO’s benefits are clearest when there is existing knowledge to preserve.
Additionally, performance improves with scale. The progression from 0.5B (mixed results) to 1.5B (strong MATH performance) to Math-1.5B (consistent gains) demonstrates that PSPO’s advantages compound as model size grows. On MATH, GR-PSPO transitions from -1.1% compared to clipping (0.5B) to +4.1pp absolute advantage (1.5B).
When fine-tuning very small base models (0.5B), the cost of discarding gradient information may be less impactful. But as models grow and accumulate knowledge, each sample becomes more costly to generate, and preserving learning signal everywhere becomes critical. PSPO’s non-zero gradients (Proposition 2.5) provide exactly this property. Unlike clipping’s hard boundary, PSPO’s behaviour-anchored smoothing provides a soft constraint that scales naturally with how far the policy deviates. Sigmoid-based alternatives such as P3O and SAPO are more vulnerable to saturation at extreme ratios and asymmetric boundaries require additional hyperparameters (SAPO and TOPR each require at least two $\tau$ values; TOPR can require up to four).
Compared to GRPO-noclip, GR-PSPO achieves similar or better performance while enabling multi-epoch training. GRPO-noclip avoids the need for trust regions by using only a single iteration over each batch, ensuring the importance ratio is always 1. This can work well when fresh on-policy samples are cheap, but generating rollouts from LLMs is computationally expensive. When reusing data, the policy diverges from the behaviour policy, making trust region mechanisms essential for stability. Additionally, Zheng et al. (2025) and NormalUhr (2025) note that larger models often require processing rollout batches in mini-batches to fit in memory, making single-iteration training impractical. PSPO enables stable multi-epoch training without sacrificing gradient information.
Refining a base model with domain knowledge is one of the most relevant for production LLM deployment. In practice, organisations fine-tune capable foundation models that already possess substantial knowledge and capabilities. The goal is improving performance on specific tasks while preserving the model’s general competence, which is exactly where our results show PSPO excels.
6 Conclusions
We introduced Probability Smoothing Policy Optimisation (PSPO) as a gradient-preserving alternative to ratio clipping in reinforcement learning for large language models, which mixes the current policy with the behaviour policy before forming the importance ratio. This blend results in a behaviour-anchored soft trust region that linearly contracts ratios around r=1r{=}1, shrinks TV/KL divergence bounds (Lemma 2.1, Corollary 2.2), and preserves non-zero gradients everywhere (Proposition 2.3).
We empirically evaluate our method by implementing PSPO within GRPO as GR-PSPO, our method demonstrates clear advantages in settings where stability matters most. On Qwen2-Math-1.5B, a model with prior supervised mathematical training, GR-PSPO outperforms baselines across all four benchmarks; giving results similar to Llama 3 8B+TOPR Roux and Bellemare (2025). On base Qwen2.5-1.5B, GR-PSPO achieves the strongest MATH performance, exceeding GRPO-clip by 4.1pp. These results support our theoretical motivation that when there is existing capability to preserve, PSPO’s behaviour-anchored smoothing prevents the destabilising updates more effectively than clipping.
Additionally, PSPO is a drop-in replacement for clipping. PSPO requires tuning of only one hyperparameter ($\alpha$, replacing $\varepsilon$), and adds no computational or memory overhead. These properties make it particularly suited to settings where data reuse is necessary and preserving existing model capabilities is critical, the common case in post-SFT reinforcement learning pipelines.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Training Details
The training for each method was run on 1 GPU. We used python version 3.10.18, and trl=0.25.1. We save checkpoint models every 1000 steps, and return the best model checkpoint at the end of training.
We used Nvidia GPUs (A100) and AMD GPUs (MI300X) for training, using only 1 GPU at a time. It took approximately 18 hours per training run for the 150M tokens, and around 10 hours for the 75M token training runs.
A.1 Tuning Method
Each method was independently tuned using a grid search over learning rate and any parameter related to the trust region (i.e. clipping range, smoothing alpha value, tau values).
We ran the initial grid search using seed 42, for 500 max global steps. The top 5 configurations were then run again on seeds 43 and 44. The best average validation reward from this was selected for the full training runs. These seeds are different to those used during training (seeds 0, 1 and 2 were used in training).
We used the learning rates =[5e07,1e06,5e06]=[5e-07,1e-06,5e-06], and the method specific values used are shown in Table 4.
Table 4: Hyperparameter grid search values for method specific parameters.
Method Trust Region Parameter Grid Search Values Notes
GRPO-clip clipping range, $\varepsilon$ [0.1, 0.2, 0.3, 0.4]
GRPO-noclip N/A N/A
P3O temperature, $\tau$ [1.0, 2.0, 4.0, 10.0] Chen et al. (2023) suggested $\tau=4.0$
SAPO temperature for positive advantages, $\tau_{positive}$ [1.0, 2.0, 3.0] Gao et al. (2025) suggested that $\tau_{negative}>\tau_{positive}$ gave the most stable training, so we only considered configurations where $\tau_{negative}\geq\tau_{positive}$.
temperature for negative advantages, $\tau_{negative}$ [1.0, 2.0, 3.0]
TOPR truncation limit for positive advantages, $b^{+}$ [1.0, 2.0, 10.0] Roux and Bellemare (2025) suggested $a^{-}=0$ and $a^{+}=1$; we searched across $b^{+}$ and $b^{-}$ values (paper uses $b^{+}=b^{-}=1$).
truncation limit for negative advantages, $b^{-}$ [1.0, 2.0, 10.0]
GR-PSPO smoothing alpha, $\alpha$ [0.1, 0.2, 0.3, 0.4]
A.2 Hyperparameter Values
A.2.1 Qwen2.5-0.5B and Qwen2.5-1.5B
For the Qwen2.5-0.5B and -1.5B models, we used the shared parameters shown in Table 5.
Table 5: Shared training hyperparameters for the 0.5B and 1.5B experiments.
Parameter Value
Effective batch size 128 (16 per device $\times$ 8 accumulation steps)
Generations per prompt 4
KL penalty ($\beta$) 0.0
Max completion length 128 tokens
Sampling temperature 0.6
Top-pp 0.85
Warmup steps 125
Precision bfloat16
The method specific parameters are shown in Table 6.
Table 6: Hyperparameters for each trust region method for Qwen2.5-0.5B and -1.5B. All methods use the DAPO loss unless otherwise noted.
Method Iterations Learning Rate Trust Region Parameters
GRPO-noclip 1 1e-6
GRPO-clip 2 1e-6 $\varepsilon=0.4$
GR-PSPO 2 1e-6 $\alpha=0.1$
P3O 2 1e-6 $\tau=4.0$
SAPO 2 1e-6 $\tau_{+}=1.0$, $\tau_{-}=3.0$
TOPR 2 1e-6 $b^{+}=1.0$, $b^{-}=1.0$
Uses GRPO-style loss to match existing implementations.
A.2.2 Qwen2-Math-1.5B
For the Qwen2-Math-1.5B model, we used the shared parameters shown in Table 5.
The method specific parameters are shown in Table 7.
Table 7: Hyperparameters for each trust region method for Qwen2-Math-1.5B. All methods use the DAPO loss unless otherwise noted.
Method Iterations Learning Rate Trust Region Parameters
GRPO-noclip 1 5e-6
GRPO-clip 2 5e-6 $\varepsilon=0.2$
GR-PSPO 2 1e-6 $\alpha=0.4$
P3O 2 1e-6 $\tau=4.0$
SAPO 2 1e-6 $\tau_{+}=1.0$, $\tau_{-}=3.0$
TOPR 2 1e-6 $b^{+}=1.0$, $b^{-}=1.0$
Uses GRPO-style loss to match existing implementations.
Appendix B Sensitivity to Smoothing Alpha
We ablate the smoothing parameter $\alpha$ on Qwen2-Math-1.5B with a learning rate of 1e-6 for 250 global steps. Performance peaks around $\alpha$=0.4, with degradation at both extremes, which could support our theory that insufficient smoothing ($\alpha$=0.1) provides weak regularisation, while excessive smoothing ($\alpha$=0.5) over-constrains updates. Table 8 shows the mean validation reward at the end of the 250 steps across different values of $\alpha$.
Table 8: Small ablation across smoothing alpha, $\alpha$, values on Qwen2-Math-1.5B.
$\alpha$ Mean Validation Reward
0.1 0.170
0.2 0.276
0.4 0.319
0.5 0.240